# Massively Multilingual Adaptation of Large Language Models Using Bilingual Translation Data

### v1

@@ 3.5 Overall Results
We evaluate the model on a comprehensive set of multilingual and bilingual benchmarks, assessing both language understanding and generation tasks. Tables 6 and 7 present the results of deterministic and generation tasks. Our EMMA-500 models are the best at machine translation (Flores200) and competitive at text classification (Taxi1500 and SIB-200) and commonsense reasoning (XCOPA and XStoryCloze). While our models are competitive in many cases, we also observe some performance degradation, such asin math tasks Performance inevitably varies across tasks and languages, reflecting trade-offs in model capacity and data representation. Appendix E presents a deeper dive into the performance details across all tasks and languages evaluated, providing insights into model behavior across multilingual datasets and benchmarks. We also provide detailed per-language results, available at https://mala-lm.github.io/emma-500-v2
@@ Acknowledgment
The work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546], and the Digital Europe Programme under grant agreement No 101195233. The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.
@@ Appendix A Frequently Asked Questions
A: Our main achievements are summarized as the contributions as written in the Section 1 Introduction. For continual pre-training, we do not modify the model architecture—we do not claim this novelty. However, our scientific novelty is more on the resource construction side, including the creation of MaLA translation corpus and continually pre-trained EMMA-500 models based on Llama 3 and 3.1 on two diverse data mixes. Besides, our empirical evaluation reveals some valuable findings that can potentially promote future development.
@@ B.1 Data Sources
Table 8 lists the corpora and collections we use as bilingual data sources in this work. A data source being bilingual means that it contains corresponding, or parallel, text records in two languages. This usually means one side has been translated, either by a machine or a human. Pedantically, massively parallel bilingual corpora can be considered a special case of multilingual corpora containing sentences in different languages that share the same meaning. In this work, some data sources have parallel text data in more than two languages, but in such cases, we always iteratively determine one as the source and one as the target language, i.e., create bilingual corpora out of them. In some source datasets, the source language is not explicitly stated, and as such we pick one language to regard as the source language without actual knowledge of the direction of translation. In such cases, we tend to pick a high-resource language, such as English, as the source language.
@@ B.2 Supported Languages
Table 9 shows the language codes of the MaLA corpus (34), where “unseen” means the languages are not used for training EMMA-500. The classification system for token counts categorizes language resources based on their size into five distinct tiers: “high” for resources exceeding 1 billion tokens, indicating a vast amount of data; “medium-high” for those with more than 100 million tokens, reflecting a substantial dataset; “medium” for resources that contain over 10 million tokens, representing a moderate size; “medium-low” for datasets with over “1 million tokens”, indicating a smaller yet significant amount of data; and finally, “low” for resources containing less than 1 million tokens, which suggests a minimal data presence. This hierarchy helps in understanding the scale and potential utility of the language resources available.
@@ B.3 Token Counts
Figure 5 shows the numbers of segments and tokens across all language pairs in the MaLA translation corpus.
@@ Appendix C Details of Data Mixes
Table 10 provides how different data types and resource categories contribute to the overall dataset composition. Our data mixes aim to make a balanced distribution over high- and low-resources. However, due to the nature of high-resource languages, they still contribute to a large portion. For example, English research papers form a substantial portion. Medium-high and medium-resource monolingual and bilingual texts contribute a lot. Medium and low-resource languages consist of 33% and 19% of bilingual and monolingual mixes, respectively.
@@ D.1 Tasks and Benchmarks
Table 11 summarizes the tasks and benchmarks used for evaluation.
@@
Text Classification (Section E.1) SIB200 (3) Accuracy 204 205 Misc
@@
Commonsense Reasoning (Section E.2) XCOPA (70) Accuracy 600 11 Misc
@@
Natural Language Inference (Section E.3) XNLI (15) Accuracy 2490 15 Misc
Machine Translation (Section E.4) FLORES-200 (16) BLEU, chrF++ 1012 204 Misc
Summarization (Section E.5) XL-Sum (29) ROUGE-L, BERTScore 2537 44 News
@@
Machine Comprehension (Section E.6) BELEBELE (10) Accuracy 900 122 Misc
@@
Math (Section E.7) MGSM direct (79) Accuracy 250 10 Misc
@@ E.2 Commonsense Reasoning
Across XCOPA and XStoryCloze evaluation results (Table 13), the four EMMA-500 variants built on Llama-3/3.1 8 B clearly lift commonsense-reasoning accuracy over their backbones. The best model, EMMA-500 Llama 3.1 8B Bi, averages 67.25% on XCOPA and 68.47% on XStoryCloze—roughly +5% over vanilla Llama-3.1 and almost ahead of, all other similar size models. The performance gains are consistent across all different languages (Tables 14 and 15). Besides, improvements concentrate in medium- and low-resource languages: EMMA-500 lifts Llama-3/3.1 by roughly +7% in medium-resource languages and flattens the usual low-resource performance drop on XStoryCloze ( 64% vs. 53% for the base), while high-resource accuracy remains competitive at around 70%.
@@ E.6 Machine Reading Comprehension
We first evaluate on the BELEBELE dataset (10), a comprehensive multilingual benchmark for machine reading comprehension, spanning 122 languages across both high- and low-resource categories. Notably, this benchmark demonstrates significant difficulty, with even the English subset posing substantial challenges to state-of-the-art language models like Llama 3. As shown in Table 23, which presents zero-shot performance across language resource groups, our continual pre-training approach on Llama 3 and 3.1, as well as LLaMAX CPT models, yields some degraded improvements over the base Llama model. However, when measured by the number of languages with improved performance, our EMMA-500 CPT is better than the base models. Contemporary models such as Aya Expanse, Llama 3.1, and Qwen 2 demonstrate more competent performance, suggesting advances in multilingual understanding and reasoning capacities. On the harder ARC-Multilingual set, the bilingual Llama-3 variant essentially ties the base model ($\approx$ 35%), while the others dip by up to one point. These results suggest that EMMA-style bilingual continual pre-training alone is insufficient for cross-language passage-question reasoning at the 8 B scale, the fundamental limitations of Llama models, and the average-based evaluation practice.
@@ E.7 Math
We use the Multilingual Grade School Math Benchmark (MGSM) (79) to evaluate the mathematical reasoning in LLMs. As shown in Table 24, we evaluate model performance using 3-shot prompting with flexible answer matching. Our assessment employs two distinct prompting strategies: (1) Direct: direct question prompting, and (2) Chain-of-Thought (CoT): prompting accompanied by step-by-step reasoning examples (97). The results show that our CPT models, as well as LLaMAX CPT models, obtain decreased average performance on direct prompting. Our EMMA-500 Llama 3.1 8B Mono model slightly improves the base model. Notably, our EMMA-500 Llama 3/3.1 CPT models significantly improve the performance on low-resource languages for both direct and CoT prompting.
Tables 25 and 26 present the per-language performance across all the languages by direct and CoT prompting.

### v2

@@ 3.5 Overall Results
We evaluate the model on a comprehensive set of multilingual and bilingual benchmarks, assessing both language understanding and generation tasks. Tables˜6 and 7 present the results of deterministic and generation tasks. Our EMMA-500 models are the best at machine translation (Flores200) and competitive at text classification (Taxi1500 and SIB-200) and commonsense reasoning (XCOPA and XStoryCloze). While our models are competitive in many cases, we also observe some performance degradation, such asin math tasks Performance inevitably varies across tasks and languages, reflecting trade-offs in model capacity and data representation. Appendix˜E presents a deeper dive into the performance details across all tasks and languages evaluated, providing insights into model behavior across multilingual datasets and benchmarks. We also provide detailed per-language results, available at https://mala-lm.github.io/emma-500-v2
@@ Acknowledgment
The work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546], and the Digital Europe Programme under grant agreement No 101195233. The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources. We thank Indraneil Paul for his contributions to preparing the code data.
@@ Appendix A Frequently Asked Questions
A: Our main achievements are summarized as the contributions as written in the Section˜1 Introduction. For continual pre-training, we do not modify the model architecture—we do not claim this novelty. However, our scientific novelty is more on the resource construction side, including the creation of MaLA translation corpus and continually pre-trained EMMA-500 models based on Llama 3 and 3.1 on two diverse data mixes. Besides, our empirical evaluation reveals some valuable findings that can potentially promote future development.
@@ B.1 Data Sources
Table˜8 lists the corpora and collections we use as bilingual data sources in this work. A data source being bilingual means that it contains corresponding, or parallel, text records in two languages. This usually means one side has been translated, either by a machine or a human. Pedantically, massively parallel bilingual corpora can be considered a special case of multilingual corpora containing sentences in different languages that share the same meaning. In this work, some data sources have parallel text data in more than two languages, but in such cases, we always iteratively determine one as the source and one as the target language, i.e., create bilingual corpora out of them. In some source datasets, the source language is not explicitly stated, and as such we pick one language to regard as the source language without actual knowledge of the direction of translation. In such cases, we tend to pick a high-resource language, such as English, as the source language.
@@ B.2 Supported Languages
Table˜9 shows the language codes of the MaLA corpus (34), where “unseen” means the languages are not used for training EMMA-500. The classification system for token counts categorizes language resources based on their size into five distinct tiers: “high” for resources exceeding 1 billion tokens, indicating a vast amount of data; “medium-high” for those with more than 100 million tokens, reflecting a substantial dataset; “medium” for resources that contain over 10 million tokens, representing a moderate size; “medium-low” for datasets with over “1 million tokens”, indicating a smaller yet significant amount of data; and finally, “low” for resources containing less than 1 million tokens, which suggests a minimal data presence. This hierarchy helps in understanding the scale and potential utility of the language resources available.
@@ B.3 Token Counts
Figure˜5 shows the numbers of segments and tokens across all language pairs in the MaLA translation corpus.
@@ Appendix C Details of Data Mixes
Table˜10 provides how different data types and resource categories contribute to the overall dataset composition. Our data mixes aim to make a balanced distribution over high- and low-resources. However, due to the nature of high-resource languages, they still contribute to a large portion. For example, English research papers form a substantial portion. Medium-high and medium-resource monolingual and bilingual texts contribute a lot. Medium and low-resource languages consist of 33% and 19% of bilingual and monolingual mixes, respectively.
@@ D.1 Tasks and Benchmarks
Table˜11 summarizes the tasks and benchmarks used for evaluation.
@@
Text Classification (Section˜E.1) SIB200 (3) Accuracy 204 205 Misc
@@
Commonsense Reasoning (Section˜E.2) XCOPA (70) Accuracy 600 11 Misc
@@
Natural Language Inference (Section˜E.3) XNLI (15) Accuracy 2490 15 Misc
Machine Translation (Section˜E.4) FLORES-200 (16) BLEU, chrF++ 1012 204 Misc
Summarization (Section˜E.5) XL-Sum (29) ROUGE-L, BERTScore 2537 44 News
@@
Machine Comprehension (Section˜E.6) BELEBELE (10) Accuracy 900 122 Misc
@@
Math (Section˜E.7) MGSM direct (79) Accuracy 250 10 Misc
@@ E.2 Commonsense Reasoning
Across XCOPA and XStoryCloze evaluation results (Table 13), the four EMMA-500 variants built on Llama-3/3.1 8 B clearly lift commonsense-reasoning accuracy over their backbones. The best model, EMMA-500 Llama 3.1 8B Bi, averages 67.25% on XCOPA and 68.47% on XStoryCloze—roughly +5% over vanilla Llama-3.1 and almost ahead of, all other similar size models. The performance gains are consistent across all different languages (Tables˜14 and 15). Besides, improvements concentrate in medium- and low-resource languages: EMMA-500 lifts Llama-3/3.1 by roughly +7% in medium-resource languages and flattens the usual low-resource performance drop on XStoryCloze ( 64% vs. 53% for the base), while high-resource accuracy remains competitive at around 70%.
@@ E.6 Machine Reading Comprehension
We first evaluate on the BELEBELE dataset (10), a comprehensive multilingual benchmark for machine reading comprehension, spanning 122 languages across both high- and low-resource categories. Notably, this benchmark demonstrates significant difficulty, with even the English subset posing substantial challenges to state-of-the-art language models like Llama 3. As shown in Table˜23, which presents zero-shot performance across language resource groups, our continual pre-training approach on Llama 3 and 3.1, as well as LLaMAX CPT models, yields some degraded improvements over the base Llama model. However, when measured by the number of languages with improved performance, our EMMA-500 CPT is better than the base models. Contemporary models such as Aya Expanse, Llama 3.1, and Qwen 2 demonstrate more competent performance, suggesting advances in multilingual understanding and reasoning capacities. On the harder ARC-Multilingual set, the bilingual Llama-3 variant essentially ties the base model ($\approx$ 35%), while the others dip by up to one point. These results suggest that EMMA-style bilingual continual pre-training alone is insufficient for cross-language passage‐question reasoning at the 8 B scale, the fundamental limitations of Llama models, and the average-based evaluation practice.
@@ E.7 Math
We use the Multilingual Grade School Math Benchmark (MGSM) (79) to evaluate the mathematical reasoning in LLMs. As shown in Table˜24, we evaluate model performance using 3-shot prompting with flexible answer matching. Our assessment employs two distinct prompting strategies: (1) Direct: direct question prompting, and (2) Chain-of-Thought (CoT): prompting accompanied by step-by-step reasoning examples (97). The results show that our CPT models, as well as LLaMAX CPT models, obtain decreased average performance on direct prompting. Our EMMA-500 Llama 3.1 8B Mono model slightly improves the base model. Notably, our EMMA-500 Llama 3/3.1 CPT models significantly improve the performance on low-resource languages for both direct and CoT prompting.
Tables˜25 and 26 present the per-language performance across all the languages by direct and CoT prompting.